shane mae and jessa can finish cleaning the house in 2 hours if it takes shane mae 3 hours than it takes jessa working alone how many hours will each girl finish the work alone
1
Expert's answer
2020-12-15T07:19:32-0500

Let X = Jessa's working time and x+3 be shane mae's.

2/x + 2/(x+3 )=1

Multiply through by x(x+3)

2(x+3) + 2x = x(x+3)

2x + 6 +2x = x+ 3x

4x + 6 = x- 3x

x- x - 6 = 0

(x-3) (x+2 )= 0

x=3 or x= -2

x=3 is the solution


Jessa can disinfect their house in 3 hours


Finding Shane Mae's working time: x+3

= 3+3

= 6

Shane Mae's can disinfect their house in 6 hours